Find the general solution. Please and thanks.

2. x(t) = 0 1 1 1 0 0 1 2(t) 1 1 0 1 3. x(t) = 1) =(t) i 1 -i
Answer #1

efem fimel amaithen arjen wecilar A B = P + x E EN G for da-1 Zz is another (A-1) = eigen verston 1 1 2 for da and 8:1 1 1 1181:3-8 Zz t + 8 1 for of 50 Sro 어니 2 Izle Zz [] then Generd salution -2 his = G -X 1 1 1 Com e 2 t 1 + 3.94 1 CScanned withazat then A = at 749 1 H2 + d3 eigen vector eigen rector corresponding to sigen value d=-1 10-1-1) 1 y Lest 1 2 Y2 I ork)td tot dt 172 then general solution G car- 201 has a G [ c1[] Gen G.S. = 2- O #(+) = 1 alt) d I 1 for eigen values A la-jI=0d = -2,-2, 2 eigen valmes finel eigen recitan corresponding eigen valios d=2 tedy, 211 22 = -2 1 1-2 03 - 271, f3y + 3O 2 -2%x (t) 1 에버 - 1 A- 11 ri 1 for eigen valises A-d2| = 0 ~ - 1-4 (1 -119.1) - Li). - 0 (2-1) (1-4) - 1-1-1-yea -2d=0 d (d-2)=0an V = -2 - 2 v, 4 Vazo ne: zuz V = ive (2.1) V, a tiva than let V₂ = 8 te, a aith V = ait then -L f V₂ 1 Lest I eigen Neetan